A company pays dividends of 5 dollars in one year. If investors require a return of 8% per annum and dividends are expected to grow at 2% what is the value of the company’ stock?
1
Expert's answer
2020-05-17T17:45:53-0400

Using Gordon Growth Model we can calculate value of company' stock:

"\frac{5}{0.08-0.02}=83 \frac{1}{3}" dollars


P.S. If you need answer in the form of a decimal fraction:

"83\frac{1}{3}\approx83.3333333"